OpenStudy (anonymous):

A rectangular prism is completely packed with 90 cubes of edge length fraction 1 over 2 inch, without any gap or overlap. Which of the following describes the volume of this rectangular prism? 11 unit cubes and 1 smaller cube of volume fraction 1 over 8 cubic inch 22 unit cubes and 4 smaller cubes of volume fraction 1 over 8 cubic inch each 11 unit cubes and 2 smaller cubes of volume fraction 1 over 8 cubic inch each 22 unit cubes and 2 smaller cubes of volume fraction 1 over 8 cubic inch each Could someone teach me the steps please? I reviewed the lesson, I still don't understand

OpenStudy (welshfella):

Hint: volume of each cube = 1/2 * 1/2 * 1/2

OpenStudy (welshfella):

oh you were given that anyway!

OpenStudy (anonymous):

so that would be 1/6

OpenStudy (welshfella):

no they are multiplied

OpenStudy (anonymous):

I'm pretty sure the volume of each cube is 1/8. Just divide 90 by 1/8 after that and you'll find the answer.

OpenStudy (welshfella):

yep

OpenStudy (anonymous):

Oh yeah, so 1/8 right? Ye

OpenStudy (anonymous):

Thank you you two, I'm supposed to be in 10th grade math buuut I have to retake 6th... Kinda embarrassing.

OpenStudy (anonymous):

It's okay! gl<3

OpenStudy (welshfella):

well multiply 90 by 1/8 and find the quotient and remainder

OpenStudy (anonymous):

While I have you two let me check my work with you if it's fine

OpenStudy (anonymous):

A-okay!

OpenStudy (anonymous):

So it would be (90*1)/8 right?

OpenStudy (anonymous):

Which would come out to be 11.25?

OpenStudy (welshfella):

No there are 11 unit cubes but how many of the smaller ones?

OpenStudy (welshfella):

90 / 8 = 11 and how many over?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Wait no, 2?

OpenStudy (welshfella):

11*8 = 88 - that equals 11 unit cubes and there are 90-88 = 2 small ones left

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so it would be c!

OpenStudy (welshfella):

right C

OpenStudy (anonymous):

Thank you for the help! I really appreciate it.

OpenStudy (welshfella):

yw